Two years ago john was 7 times as old as his son but in 3 years he wil be 4 times as old as the boy. How old is john?
Answer by rothauserc(4718)

Let x be the age of John and y be the age of his son
:
1) x-2 = 7 * (y-2)
2) x+3 = 4 * (y+3)
:
Use equation 1 and solve for x
:
x-2 = 7y-14
x = 7y-12
:
Substitute for x in equation 2
:
7y-12+3 = 4y+12
3y = 21
:
y = 7
x = 49-12 = 37
:
*************************
John is 37 years old
John's son is 7 years old
*************************
: